OpenStudy (anonymous):

P = 2l + 2w + 2k for k

OpenStudy (anonymous):

P=2l +2w +2k p-2l=2w+2k p-2l-2w=2k (p/2)-l-w=k

OpenStudy (anonymous):

\[P = 2(l+w+k)\] \[P/2=l+w+k\] \[k = (P/2)-l-w\]